An aircraft traveling with a velocity relative to the air of 320 km/h [S62°W] passes over Winnipeg.The wind velocity is 72 km/h [S].Determine the velocity of the plane relative to the ground.


1
Expert's answer
2022-03-10T10:18:22-0500

"u=v+v'\cos 62°=354~\frac{km}h."